OpenStudy (anonymous):

Given the trinomial 5x2 - 2x - 3, predict the type of solutions. Two rational solutions One rational solution Two irrational solutions Two complex solutions

OpenStudy (triciaal):

Descates rule of signs + - - yes no one positive and one negative root

OpenStudy (triciaal):

5x^2 -2x -3 (5x +3)(x-1) x = -3/5; x = 1

OpenStudy (anonymous):

ahhh ok i got it

OpenStudy (triciaal):

the degree of the polynomial will also determine the number of roots 2nd degree 2 roots 3rd 3 roots 4th degree 4 roots etc

OpenStudy (anonymous):

so its a

OpenStudy (triciaal):

i agree

OpenStudy (anonymous):

thx

OpenStudy (triciaal):

you are welcome

OpenStudy (mathmath333):

u can also use discremenant

OpenStudy (mathmath333):

\(\tt b^2 - 4ac > 0 .... two ~~real ~~roots \\ b^2 - 4ac = 0 .... one ~~real ~~root \\ b^2 - 4ac < 0 .... two ~~complex~~ roots\)

OpenStudy (triciaal):

yes you can also use the discriminant where a, band c are the coefficients of a quadratic equation